OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - 6x + 8? (4 points) x = -4, 2 x = -4, -2 x = 4, 2 x = 4, -2

OpenStudy (anonymous):

help!!

OpenStudy (sleepyhead314):

Hello! sorry for the kinda slow reply .-. do you know how to factor? aka backwards FOILing?

OpenStudy (anonymous):

yes i know how to foil first outers inners last

OpenStudy (sleepyhead314):

that's good :) but now we have to do it backwards for example x^2 - 8x + 12 = (x)(x) - (6x) - (2x) + (6)(2) = (x - 6)(x - 2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you walk me through it please?

OpenStudy (sleepyhead314):

of course :) first we have to understand that in our equation: \( x^2-6x+8 \) we have to find two numbers (x + #)(x + #) that will go here ^ ^ these two numbers should \(multiply~up~to~positive~8\) and \(add~up~to~negative~6\)

OpenStudy (sleepyhead314):

so let's first list the pairs of numbers that can multiply up to a positive 8 :)

OpenStudy (anonymous):

ok

OpenStudy (sleepyhead314):

this would include \( 1~and~8 \) \( -1~and~-8\) etc can you find a pair that adds up to \(-6\) ? [adds up like \(-1 + -8 = -9\) ]

OpenStudy (anonymous):

ok i see

OpenStudy (sleepyhead314):

1 and 8 -1 and -8 2 and 4 -2 and -4 1+8=9 -1 + -8 = -9 2+4= 6 -2 + -4 = -6 <-- so the pair of -2 and -4 works

OpenStudy (sleepyhead314):

all we have to do now is to plug "-2" and "-4" into (x + __)(x + __) ^ ^ to have our factored version of the equation

OpenStudy (sleepyhead314):

\(Warning!\) Don't select an answer yet!

OpenStudy (sleepyhead314):

(x - 2)(x - 4) now the "zeros" of this would be when each part is set equal to "zero" x - 2 = 0 [solve for x] x - 4 = 0 [solve for x] noww you will get your answers :)